# Rigidity of holomorphic maps between bounded symmetric domains preserving Shilov boundary

### v1

@@ Abstract
We introduce the concept of orthogonal structure on complex Grassmannians. Based on this structure, we define the notion of orthogonal mappings. This class of maps generalizes holomorphic maps between type I bounded symmetric domains that preserve Shilov boundary. By analyzing the geometric properties of these orthogonal mappings, we obtain rigidity results for such mappings. As an application, we establish a rigidity result for holomorphic maps from rank $1$ type I bounded symmetric domains $\Omega_{1,s}$ (complex balls $B^{s}$) to higher-rank domains $\Omega_{r^{\prime},s^{\prime}}$ that preserve their corresponding Shilov boundaries, where $s\geq 2$ and $2\leq r^{\prime}\leq s^{\prime}$. More specifically, we show that (1): such maps are constant when $s^{\prime}-r^{\prime}<s-1$; (2) for $s-1\leq s^{\prime}-r^{\prime}<2s-2$, these maps reduce to standard linear embeddings after normalization by applying automorphisms on both sides. Our results are optimal and generalize the well-known optimal bounds for proper mappings between rank $1$ cases and CR maps between Shilov boundaries of higher-rank type I bounded symmetric domains.
@@ 1 Introduction
The rigidity for the proper holomorphic maps between balls (rank 1 bounded symmetric domains) has received considerable attention with its origins tracing back to Poincaré’s work for $2$-dimensional unit ball. For unit balls in different dimensions, this field has been extensively explored by many mathematicians from various directions including algebraic geometry, Cauchy-Riemann geometry, and partial differential equation, etc. with major contributions from Cima-Suffridge [CS], D’Angelo [Da1], [Da2], D’Angelo-Kos-Riehl [DKR], D’Angelo-Lebl [DL1],[DL2], Faran [Fa], Forstneric [F1],[F2], Huang [Hu1],[Hu2], Huang-Ji [HJ], and Huang-Ji-Yin [HJY], among others.
@@
Let $r\leq s\in\mathbb{N}$. Type I bounded symmetric domains $\Omega_{r,s}=\{Z\in\mathbb{C}^{r\times s}\mid I_{r}-ZZ^{H}>0\}$ with rank $r$, along with their Shilov boundaries $S(\Omega_{r,s})=\{Z\in\mathbb{C}^{r\times s}\mid I_{r}-ZZ^{H}=0\}$ where $Z^{H}=\bar{Z}^{t}$, have been studied by Kim-Zaitsev in [KZ13] with regard to CR rigidity. The Shilov boundary imposes stringent constraints on holomorphic maps, forcing them to align with canonical substructures. Kim-Zaitsev show that CR embeddings between $S(\Omega_{r,s})$ and $S(\Omega_{r^{\prime},s^{\prime}})$ are linear under the assumptions $s^{\prime}-r^{\prime}<2(s-r)$, $2\leq r<s$, $2\leq r^{\prime}<s^{\prime}$, and certain smoothness condition, using the Cartan moving frame method. In a very recent paper [Kim2], the author also describes explicitly the holomorphic mappings between Type I bounded symmetric domains that preserve Shilov boundary. These results crucially depend on the source domain having $r$, $r^{\prime}\geq 2$. The analog of Kim-Zaitsev’s theorem in the rank $1$ case was solved earlier by Faran [Fa].
@@
We introduce a novel geometric approach to this problem by defining an orthogonal structure on Grassmannians and studying associated orthogonal mappings between them. This framework, integrating orthogonality with an effective analysis of linear subspaces and projective geometry techniques, yields rigidity results without resorting to heavy machinery. These orthogonal mappings slightly generalize holomorphic maps between Type I bounded symmetric domains that preserve their Shilov boundaries. Consequently, our approach provides new rigidity results for holomorphic mappings from rank $1$ bounded symmetric domains to higher-rank domains preserving the Shilov boundary.
@@
where $z=(z_{1},\ldots,z_{r+s})$ and $w=(w_{1},\ldots,w_{r+s})$.
@@
Consider the Grassmannian variety $G(r,r+s)$, whose points bijectively correspond to $r$-dimensional linear subspace in $\mathbb{C}^{r+s}$. Let $p\in G(r,r+s)$. We use $V_{p}$ to denote the $r$-dimension linear subspace corresponding to $p$. An orthogonal structure can be defined on $G(r,r+s)$ induced by the inner product $\langle\cdot,\cdot\rangle_{r,s}$. We call two points $p$, $q$ orthogonal in $G(r,r+s)$ if and only if $V_{p}\perp V_{q}$ in $\mathbb{C}^{r,s}$, denoted by $p\perp q$. We then denote by $\mathcal{G}(r,s)$ the Grassmannian variety $G(r,r+s)$ equipped with the orthogonal structure. This leads to the definition of local orthogonal maps (Definition 2.2). From Remark after Definition 2.2, we shall see that the local orthogonal maps from $\mathcal{G}(r,s)$ to $\mathcal{G}(r^{\prime},s^{\prime})$ and the local holomorphic map between Shilov boundary $S(\Omega_{r,s})$ and $S(\Omega_{r^{\prime},s^{\prime}})$ are actually almost the same set of mappings although the former could be defined slightly more generally. Moreover, that a holomorphic map $F$ is orthogonal is essentially an alternative way to express that $F$ preserves the Shilov boundary, both of which come from the standard polarization argument on the function equations. Hence we can combine orthogonality with an effective analysis of linear subspaces and techniques in projective geometry to obtain rigidity theorems for holomorphic mappings between Shilov boundaries. We are now ready to state our main theorems. Further explanation and details can be found in Section 3.
In this paper, we prove:
@@
Let $s\geq 2$, $2\leq r^{\prime}\leq s^{\prime}\in\mathbb{N}$ and let $F$ be a holomorphic map from the Type I bounded symmetric domain $\Omega_{1,s}$ to $\Omega_{r^{\prime},s^{\prime}}$ that preserves the Shilov boundary.
@@
    $\bf z\to\left(\begin{array}[]{ccc}I&0&0\\ 0&\bf z&0\\ \end{array}\right).$
We note that our proof of Theorem 3.6 is centered around the orthogonality preservation property and the dimensional analysis deduced from it. This new perspective enables us to avoid sophisticated computations as in the previous works, and makes our paper significantly shorter.
@@ 2 Orthogonal structure on Grassmannian
Let $r$, $s$ be positive integers and $r\leq s$. Consider the Grassmannian variety $G(r,r+s)$, whose closed points are in bijection with $r$-dimensional vector subspaces $V\subset\mathbb{C}^{r+s}$. Fix the standard basis $(e_{1},\ldots,e_{r+s})$ of $\mathbb{C}^{r+s}$. A subspace $V$ can be represented as the row space of an $r\times(r+s)$ matrix:
@@
Let $U\subseteq\mathcal{G}(r,s)$ be a connected open set containing a null point. A holomorphic map $F:U\to\mathcal{G}(r^{\prime},s^{\prime})$ is called orthogonal if $F(p)\perp F(q)$ for all $p,q\in U$. We also simply call $F$ a local orthogonal map from $\mathcal{G}(r,s)$ to $\mathcal{G}(r^{\prime},s^{\prime})$.
Remark: For the sake of simplicity, in our definition we require the domain $U$ to contain a null point, i.e. to have a non-empty intersection with $S(\Omega_{r,s})$ so that the orthogonality or sign-preserving condition will not become vacuous. (Otherwise, it could happen that there are no orthogonal points in $U$.) With this requirement, Proposition 2.4 and Proposition 2.5 below will imply that orthogonal maps and holomorphic maps between Shilov boundaries are the same set of mappings. However, we remark that in principle one can define orthogonal maps more generally without requiring $U\cap\Omega_{r,s}=\emptyset$ (but then we need to assume that $U$ contains at least a pair of orthogonal points) and some of our results (e.g. Theorem 3.6) will still be valid. Thus, we may say that a local orthogonal map is a slightly more general object.
@@
Every local orthogonal map $F:\mathcal{G}(r,s)\to\mathcal{G}(r^{\prime},s^{\prime})$ preserves null points, i.e., $F(S(\Omega_{r,s}))\subseteq S(\Omega_{r^{\prime},s^{\prime}})$.
@@
For $\mathcal{Z}=[I_{r},Z]\in\mathcal{G}(r,s)$, a point $\mathcal{W}=[I_{r},W]$ lies in $\mathcal{Z}^{\perp}\cap U^{\prime}$ precisely when $h_{kl}(Z,\bar{W})=0$ for all $k,l$. By (1), this implies $F(\mathcal{W})\in F(\mathcal{Z})^{\perp}$. Hence, $F$ preserves orthogonality locally, proving $F|_{U^{\prime}}$ is an orthogonal map.
@@ 3 Rigidity of local orthogonal maps between Grassmannians
Let $s>1$, $1\leq r^{\prime}\leq s^{\prime}\in\mathbb{N}$, and let $F$ be a local orthogonal map from $U\subset\mathcal{G}(1,s)=\mathbb{P}^{1,s}$ to $\mathcal{G}(r^{\prime},s^{\prime})$. If
@@
Let $s\geq 2$, $2\leq r^{\prime}\leq s^{\prime}\in\mathbb{N}$ and $F$ be a holomorphic map from Type I bounded symmetric domain $\Omega_{1,s}$ to $\Omega_{r^{\prime},s^{\prime}}$ preserving the Shilov boundaries.
@@
Acknowledgements The author extends her gratitude to Prof. Sui-Chung Ng for carefully reading the first version of this article, and Prof. Xiaojun Huang and Prof. Chenyang Xu for their insightful comments and valuable suggestions on this work. This research was supported by the National Natural Science Foundation of China (NSFC) under Grant Nos. 12471042 and 12471078.

### v2

# Rigidity of holomorphic maps between Shilov boundaries of type-I bounded symmetric domains

@@ Abstract
We introduce the concept of orthogonal structure on complex Grassmannians. Based on this structure, we define the notion of orthogonal mappings. This class of maps generalizes holomorphic maps between the Shilov boundaries of type-I bounded symmetric domains. By analyzing the geometric properties of these orthogonal mappings, we obtain rigidity results for such mappings. As an application, we establish a rigidity result for the holomorphic maps from the Shilov boundary of rank $1$ type I bounded symmetric domain $\Omega_{1,s}$ (i.e. unit spheres) to the Shilov boundary of a higher-rank type-I domain $\Omega_{r^{\prime},s^{\prime}}$, where $s\geq 2$ and $2\leq r^{\prime}\leq s^{\prime}$. More specifically, we show that: (1) such maps are constant when $s^{\prime}-r^{\prime}<s-1$; (2) for $s-1\leq s^{\prime}-r^{\prime}<2s-2$, they reduce to the standard linear embeddings after normalization by applying automorphisms on both sides. Our results are optimal and generalize the well-known optimal bounds for proper mappings between rank $1$ cases and CR maps between Shilov boundaries of higher-rank type I bounded symmetric domains.
@@ 1 Introduction
The rigidity for the proper holomorphic maps between balls (rank 1 bounded symmetric domains) has received considerable attention with its origins tracing back to Poincaré’s work for $2$-dimensional unit ball. For unit balls in different dimensions, this field has been extensively explored by many mathematicians from various directions including algebraic geometry, Cauchy-Riemann geometry, and partial differential equation, etc. see for example Cima-Suffridge [CS], D’Angelo [Da1], [Da2], D’Angelo-Kos-Riehl [DKR], D’Angelo-Lebl [DL1],[DL2], Faran [Fa], Forstneric [F1],[F2], Huang [Hu1],[Hu2], Huang-Ji [HJ], and Huang-Ji-Yin [HJY], among others.
@@
Let $r\leq s\in\mathbb{N}$. Type I bounded symmetric domains $\Omega_{r,s}=\{Z\in\mathbb{C}^{r\times s}\mid I_{r}-ZZ^{H}>0\}$ with rank $r$, along with their Shilov boundaries $S(\Omega_{r,s})=\{Z\in\mathbb{C}^{r\times s}\mid I_{r}-ZZ^{H}=0\}$ where $Z^{H}=\bar{Z}^{t}$, have been studied by Kim-Zaitsev in [KZ13] with regard to CR rigidity. The Shilov boundary imposes stringent constraints on holomorphic maps, forcing them to align with canonical substructures. Kim-Zaitsev show that CR mappings between $S(\Omega_{r,s})$ and $S(\Omega_{r^{\prime},s^{\prime}})$ are linear under the assumptions $s^{\prime}-r^{\prime}<2(s-r)$, $2\leq r<s$, $2\leq r^{\prime}<s^{\prime}$, and certain smoothness condition, using the Cartan moving frame method. In a very recent paper [Kim2], the author also describes explicitly the holomorphic mappings between Type I bounded symmetric domains that preserve Shilov boundary. These results crucially depend on the domains having $r$, $r^{\prime}\geq 2$. The analog of Kim-Zaitsev’s theorem in the rank $1$ case was solved earlier by Faran [Fa].
@@
We introduce a novel geometric approach to this problem by defining an orthogonal structure on Grassmannians and studying associated orthogonal mappings between them. This framework, integrating orthogonality with an effective analysis of linear subspaces and projective geometry techniques, yields rigidity results without resorting to heavy machinery. These orthogonal mappings slightly generalize the holomorphic maps between the Shilov boundaries of Type I bounded symmetric domains. Consequently, our approach provides new rigidity results for such maps. Here, by a holomorphic map between the Shilov boundaries, we mean a holomorphic map $F$ defined on a neighborhood $U$ of a point on $S(\Omega_{r,s})$ such that $F(U\cap S(\Omega_{r,s}))\subset S(\Omega_{r^{\prime},s^{\prime}})$.
@@
where $z=(z_{1},\ldots,z_{r+s})$ and $w=(w_{1},\ldots,w_{r+s})$. Denote $\mathbb{P}\mathbb{C}^{r,s}=\mathbb{P}^{r,s}$.
@@
Consider the Grassmannian variety $G(r,r+s)$, whose points bijectively correspond to the $r$-dimensional linear subspaces in $\mathbb{C}^{r+s}$. Let $p\in G(r,r+s)$. We use $V_{p}$ to denote the $r$-dimension linear subspace corresponding to $p$. An orthogonal structure can be defined on $G(r,r+s)$ induced by the inner product $\langle\cdot,\cdot\rangle_{r,s}$. We call two points $p$, $q$ orthogonal in $G(r,r+s)$ if and only if $V_{p}\perp V_{q}$ in $\mathbb{C}^{r,s}$, denoted by $p\perp q$. We then denote by $\mathcal{G}(r,s)$ the Grassmannian variety $G(r,r+s)$ equipped with the orthogonal structure. This leads to the definition of local orthogonal maps.
Definition 1.1.
Let $U\subseteq\mathcal{G}(r,s)$ be a connected open set containing a null point. A holomorphic map $F:U\to\mathcal{G}(r^{\prime},s^{\prime})$ is called orthogonal if $F(p)\perp F(q)$ for all $p,q\in U$ such that $p\perp q$. We also simply call $F$ a local orthogonal map from $\mathcal{G}(r,s)$ to $\mathcal{G}(r^{\prime},s^{\prime})$.
In our definition of orthogonal maps, we required the domain U to contain a null point, ensuring the orthogonality condition would not be vacuous. However, this condition is not essential to the definition. Orthogonal maps can be defined more generally, requiring only that $U$ contains at least one pair of orthogonal points. From Proposition 2.5, we shall see that the local orthogonal maps from $\mathcal{G}(r,s)$ to $\mathcal{G}(r^{\prime},s^{\prime})$ and the local holomorphic maps between Shilov boundary $S(\Omega_{r,s})$ and $S(\Omega_{r^{\prime},s^{\prime}})$ are actually almost the same set of mappings although the former could be defined slightly more generally.
We are now ready to state our main theorems. Further explanation and details can be found in Section 3.
@@
Let $s\geq 2$, $2\leq r^{\prime}\leq s^{\prime}\in\mathbb{N}$, and let $U\subset\mathbb{C}^{s}$ be an open neighborhood of a point in $S(\Omega_{1,s})$. let $F$ be a holomorphic map from $U$ to $\mathbb{C}^{r^{\prime}s^{\prime}}$ satisfying $F(U\cap S(\Omega_{1,s}))\subset S(\Omega_{r^{\prime},s^{\prime}})$.
@@
    $z\to\left(\begin{array}[]{ccc}I&0&0\\ 0&z&0\\ \end{array}\right).$
We note that our proof of Theorem 1.2 is centered around the orthogonality preservation property and the dimensional analysis deduced from it. This new perspective enables us to avoid sophisticated computations as in the previous works, and makes our arguments considerably shorter.
@@ 2 Orthogonal structure on Grassmannian
Let $r$, $s$ be positive integers and $r\leq s$. Consider the Grassmannian variety $G(r,r+s)$, whose closed points are in bijection with $r$-dimensional vector subspaces $V\subset\mathbb{C}^{r+s}$. Fix the standard basis $(e_{1},\ldots,e_{r+s})$ of $\mathbb{C}^{r+s}$. A linear subspace $V$ can be represented as the row space of an $r\times(r+s)$ matrix:
@@
Let $U\subseteq\mathcal{G}(r,s)$ be a connected open set containing a null point. A holomorphic map $F:U\to\mathcal{G}(r^{\prime},s^{\prime})$ is called orthogonal if $F(p)\perp F(q)$ for all $p,q\in U$ such that $p\perp q$. We also simply call $F$ a local orthogonal map from $\mathcal{G}(r,s)$ to $\mathcal{G}(r^{\prime},s^{\prime})$.
Remark: For the sake of simplicity, in our definition we require the domain $U$ to contain a null point, i.e. to have a non-empty intersection with $S(\Omega_{r,s})$ so that the orthogonality condition will not become vacuous. (Otherwise, it could happen that there are no orthogonal points in $U$.) With this requirement, Proposition 2.4 and Proposition 2.5 below will imply that orthogonal maps and holomorphic maps between Shilov boundaries are the same set of mappings. However, we remark that in principle one can define orthogonal maps more generally without requiring $U\cap S(\Omega_{r,s})\neq\emptyset$ (but then we need to assume that $U$ contains at least a pair of orthogonal points) and some of our results (e.g. Theorem 1.2) will still be valid. Thus, we may say that a local orthogonal map is a slightly more general object.
@@
Every local orthogonal map $F:\mathcal{G}(r,s)\to\mathcal{G}(r^{\prime},s^{\prime})$ preserves null points.
@@
For any point $\mathcal{Z}=[I_{r},Z]\in\mathcal{G}(r,s)$, a point $\mathcal{W}=[I_{r},W]$ lies in $\mathcal{Z}^{\perp}\cap U^{\prime}$ precisely when $h_{kl}(Z,\bar{W})=0$ for all $k,l$. By (1), this implies $F(\mathcal{W})\in F(\mathcal{Z})^{\perp}$. Hence, $F$ preserves orthogonality locally, proving $F|_{U^{\prime}}$ is an orthogonal map.
@@ 3 Rigidity of local orthogonal maps between Grassmannians
Let $s\geq 2$, $1\leq r^{\prime}\leq s^{\prime}\in\mathbb{N}$, and let $F$ be a local orthogonal map from $U\subset\mathcal{G}(1,s)=\mathbb{P}^{1,s}$ to $\mathcal{G}(r^{\prime},s^{\prime})$. If
@@
Let $s\geq 2$, $2\leq r^{\prime}\leq s^{\prime}\in\mathbb{N}$, and let $U\subset\mathbb{C}^{s}$ be an open neighborhood of a point in $S(\Omega_{1,s})$. let $F$ be a holomorphic map from $U$ to $\mathbb{C}^{r^{\prime}s^{\prime}}$ satisfying $F(U\cap S(\Omega_{1,s}))\subset S(\Omega_{r^{\prime},s^{\prime}})$.
@@
Acknowledgements The author is deeply grateful to Prof. Sui-Chung Ng for the invaluable discussions, constructive suggestions and detailed feedback throughout the development of this paper. She also wishes to thank Prof. Zaitsev whose comments not only has led to better phrasing of Corollary 1.3 but also helped refining other parts of the work. Additionally, thanks are due to Prof. Xiaojun Huang and Prof. Chenyang Xu for their insightful comments and valuable suggestions. This research was supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12471042 and 12471078.